Owl Shift, Wednesday, August 28, 2002 00:00 Shift begins. Expert: A. Gasparian. Worker: E. Clinton 03:00 MCC starts to deliver beam to Hall B. Beam inconsistent. 03:30 4 nA of CW beam from MCC. 03:40 Beam down 05:05 4 nA of CW beam 05:35 Harp scans started. Tagger scans-no radiator and radiator B Photon scans-no primakoff target and 10^-3 06:30 Scans done. 06:45 Determined target ladder horizontal centers are roughly -2.8 millimeters from nominal values. Reason: Nominl values used blue prints and limit switch placements to determine position. No effort was made to determine if beam is off from nominal "blue print" center.